# From strong to weak correlations in breathing-mode kagome van der Waals materials: Nb3(F,Cl,Br,I)8 as a robust and versatile platform for many-body engineering

### v1

@@ II.2.2 Molecular Orbital Basis
Having established the breathing mode distortion as the mechanism producing the disentangled flat bands, we are now ready to study the correlation trend across Nb3X8 family. In Fig. 3, we show the bulk DFT band structures in the LT phases. As expected, we find two relatively flat bands around the Fermi level in all cases. In the case of Nb3F8, both bands cross the Fermi level, while for the others, we consistently find one fully occupied and one completely unoccupied band. For in-plane momenta ($k_{z}=0$), the two bands are clearly split apart, while for the out-of-plane momentum direction $\Gamma-A$, we observe a stronger dispersion, and the gap nearly closes. As will become clear from the subsequent Wannierization discussed below, this $k_{z}$ behaviour is a result of the interlayer dimerization of trimers within the strongly hybridized bilayers. Thus, in $k_{z}$, the low-energetic electronic structure is reminiscent of the behaviour of the Su-Schrieffer-Heeger (SSH) model [44], which consequently yields a gap in this direction, due to the mismatch of the alternating out-of-plane hopping matrix elements. From Cl to Br and I, we already see the first clear trend in the material class: the in-plane gap between these two bands steadily increases.
@@
Figure 4: Wannier Molecular Orbitals of Nb3Br8. $\psi_{1}$ and $\psi_{2}$, along with their bonding ($+$) and antibonding (--) combinations, $\psi^{(\pm)}=(\psi_{1}\pm\psi_{2})/\sqrt{2}$.
@@
To prepare for the downfolding and to further analyze the low-energetic band structure, we construct maximally localized Wannier functions for the two low-energy bands. In all four cases, we find that the bands can be perfectly Wannier-interpolated within a basis spanned by two molecular orbitals $\psi_{1}$ and $\psi_{2}$, which are centered on trimers of adjacent layers and which are directly above each other, see Fig. 4. In this molecular orbital basis, we find similar contributions of both orbitals to both low-energy bands. However, upon diagonalization of the single-particle Wannier Hamiltonian, we find that the lower (occupied) band is well described by a bonding orbital $\psi^{(+)}$ formed by the direct sum of the trimer-centered molecular orbitals $\psi^{(+)}=(\psi_{1}+\psi_{2})/\sqrt{2}$, while the upper (unoccupied) band is well described by the corresponding anti-bonding orbital $\psi^{(-)}=(\psi_{1}-\psi_{2})/\sqrt{2}$, as illustrated in Fig. 4. For the subsequent analysis, it will be helpful to switch back and forth between the molecular orbital basis ($\psi_{i}$) and the bonding/anti-bonding basis ($\psi^{(\pm)}$).
@@
Upon projecting the Kohn-Sham states to the $\psi^{(\pm)}$ basis, we see that this interpretation holds especially for in-plane momenta. For out-of-plane $k_{z}$ momenta, however, we observe that $\psi^{(\pm)}$ states mix, which is a result of the small, but finite wavefunction overlap between the dimerized layers. This admixture thus verifies the SSH-like behaviour in the $k_{z}$ direction. This is consistent with Ref. 45 within which Nb3Br8 and Nb3I8 are classified as obstructed atomic insulators  [^1] See https://www.topologicalquantumchemistry.fr/#/detail/25767 and https://www.topologicalquantumchemistry.fr/#/detail/421609.. Although the Cl, Br, and I compounds all share the classification as obstructed atomic insulators, the details and specifically the size of the DFT gap change from material to material, as also shown by the Wannier hopping parameters in Fig. 5 and as discussed in more detail in the following.
@@
II.3 Downfolded Many-Body Hamiltonians from Constrained Random Phase Approximation Calculations
Figure 5: Molecular Orbital Hubbard Model Parameters of Nb3X8. (a) In-plane $t_{\parallel}$ and out-of-plane hopping matrix elements elements $t_{\perp}^{(s)}$ and $t_{\perp}^{(w)}$ within the molecular orbital basis. (b) cRPA Coulomb interaction matrix elements: on-site $U_{0}$, first nearest in-plane neighbors $U_{\parallel}$, and first and second nearest out-of-plane neighbors, $U_{\perp}^{(s)}$, and $U_{\perp}^{(w)}$, respectively, within the molecular orbital basis. The light green marker in panel (a) indicates varying in-plane hopping values for Nb3Cl8 due to reduced symmetry.
@@ II.3 Downfolded Many-Body Hamiltonians from Constrained Random Phase Approximation Calculations
for all compounds. Here $c_{i}^{\dagger}$ and $c_{j}$ are electron creation and annihilation operators in molecular orbitals at sites $i$ and $j$, respectively, $t_{ij}$ are single-particle (hopping) and $U_{ijkl}$ partially screened two-particle (Coulomb) matrix elements. The latter are evaluated within the constrained random phase approximation (cRPA) [47], which accounts for all RPA-level screening processes except those taking place within the molecular orbitals $\psi_{i}$. In Fig. 5, we show a subset of these matrix elements focusing on the local and nearest-neighbor density-density Coulomb matrix elements ($U_{ijji}$) as well as on intra- and inter-layer hopping matrix elements. For a complete overview, see Appendix .1.
@@
We find that the in-plane nearest-neighbor hopping ($t_{\parallel}$) is relatively small in all compounds ($\lesssim 25\,$meV in magnitude, see green markers in Fig. 5 (a), which reflects the similar rather flat in-plane (constant $k_{z}$) dispersion with Nb3Br8 having the smallest $t_{\parallel}$, and thus flattest in-plane dispersion. The strong out-of-plane nearest-neighbor hopping ($t^{(s)}_{\perp}$), in turn, grows significantly in magnitude from Cl ($-127\,$meV) to Br ($-162\,$meV) and to I ($-208\,$meV), see red markers in Fig. 5a. The “weak link” in the out-of-plane directions ($t^{(w)}_{\perp}$) is vanishingly small for Cl ($-3\,$meV) and only mildly growths towards I ($-24\,$meV), see blue markers in Fig. 5. This reflects the increasing bilayer hybridization between the inter-layer dimerized trimers, leading to the growth of the bonding/anti-bonding single particle splitting, as observed around $\Gamma$ in Fig. 3. In Nb3F8*, the in- and out-of-plane hoppings are of similar amplitude, such that there is no opening of a full dimerization gap throughout the Brillouin zone and only a small splitting at $\Gamma$ is observed. To analyze the gap at the high symmetry point $A$, we need to take into account that there are three adjacent layer neighbors in the “weak-link” directions, such that the SSH gap of a nearest-neighbor model would be given by $2\left|t^{(s)}_{\perp}-\sqrt{3}t^{(w)}_{\perp}\right|$. This increases from about $243$ to $332$ meV going from Cl to I, indicating an increase of the dimerization strength (in Nb3F8*, it is only 444\,4meV). The DFT gaps at $A$ (where the SSH gap opens) are, however smaller in Br and I as compared to Cl. We understand this as a result of long-range hoppings in the out-of-plane direction, which are indeed (slightly) more pronounced in Br and I than in Cl and F.
@@
The partially screened Coulomb interactions are dominated by density-density interactions. This is a result of having just a single molecular orbital being centered in each of the adjacent, strongly hybridized monolayers. Non-density-density interactions, such as Hund’s exchange interactions, are thus non-local, which significantly suppresses their strengths (to a few meV). Due to reduced screening in the layered vdW structures, the density-density Coulomb interactions are also non-local and show a significant long-ranged tail. The latter is visible in Fig. 5 (b) by comparing the local interactions $U_{0}$ to the nearest-neighbor ones and becomes clear by comparing the full non-localities as further detailed in Appendix .1. Overall, the density-density interactions show a consistent decreasing trend from F to I. This results from two effects. First, the molecular orbital spread increases from F to I from 4.44.44.4~{}4.4Å2 to 14.414.414.4~{}14.4Å2, which yields a decrease of the bare interactions $v_{ijkl}$. Simultaneously, the effective dielectric screening $\varepsilon$ increases from F to I, since the other bands (especially Nb $t^{2}_{2g}$ and Nb $e_{g}$) move closer to the Fermi level (see Fig. 3). Taken together, this yields the observed decrease in $U_{ijkl}$. In contrast to the out-of-plane hopping matrix elements, we do not see a significant bilayer hybridization effect in the non-local density-density Coulomb interactions. Comparing $U^{(w)}_{1,\perp}$ and $U^{(s)}_{1,\perp}$ in Fig. 5 (b) (red and blue markers), we find only small differences, which can be explained by the slightly different distances between the molecular orbital centers in these two directions.
@@
We further see that the local Coulomb interactions $U_{0}$ are very large in comparison to all hopping terms with $U_{0}\approx 2.4\,$eV in Nb3F8*, 1.51.51.5\,1.5eV in Nb3Cl8, 1.251.251.25\,1.25eV in Nb3Br8, and 0.80.80.8\,0.8eV in Nb3I8. Thus, the $U_{0}/t_{\parallel}$ ratios are always larger than $80$, while the $U_{0}/t_{\perp}$ ratio decreases from about $12$ in Nb3Cl8 to around $4$ in Nb3I8. These ratios point towards possible strong Coulomb-interaction induced correlation effects, which might change from system to system. To investigate this and quantify the correlation strengths, we proceed with solving the generalized Hubbard models using cluster dynamical mean-field theory.
@@
II.4 Cluster Dynamical Mean Field Theory Self-energies at Half-Filling
@@
Figure 6: Nb3X8 Matsubara Self-Energies. (a) Real and (b) imaginary part of the self energy on Matsubara frequencies in the bonding/anti-bonding basis. Markers represent Hubbard-I results, lines represent cluster DMFT results for Nb3Cl8, Nb3Br8 & Nb3I8.
@@
Using our cRPA-based downfolded molecular-orbital models, we calculate the self-energies at half-filling within dynamical mean-field theory (DMFT) and its Hubbard-I approximation [48, 49]. Specifically, we define the impurity problem as the two inter-layer dimerized molecular (trimer) orbitals within each bilayer and include the local $U_{0}$, the strong inter-layer hybridization $t^{(s)}_{\perp}$ as well as all inter-layer Coulomb interaction elements $U_{ijkl}$ (among which $U^{(s)}_{\perp}$ is by far the largest, see Appendix .1). Accordingly, we deal with a cluster DMFT problem [50, 51, 19]. Fig. 6 shows the resulting Matsubara frequency self-energies in the diagonal $\psi^{(\pm)}$ basis for all compounds. In this basis, the imaginary parts of the self-energy are the same for both channels ($\mathrm{Im}\Sigma_{(+)}=\mathrm{Im}\Sigma_{(-)}$), while the real parts differ only in their signs ($\mathrm{Re}\Sigma_{(+)}=-\mathrm{Re}\Sigma_{(-)}$ in the case of Cl, Br, and I).
@@
The real part, as depicted in Fig. 6 (a), shows that correlation effects enhance the bonding/anti-bonding splitting, similar to the Stoner enhancement of the exchange splitting in magnetic systems. In I, the real part is almost frequency-independent, while it becomes strongly frequency-dependent moving towards Cl. This is a sign of the increasing importance of dynamic correlation effects, which necessitate (cluster) DMFT calculations. In the imaginary part depicted in Fig 6 (b), we again find a clear trend: the overall amplitudes of the self-energy gradually increase from I to Br, Cl, and F, which is a sign of the increasing degree of correlation.
@@
II.5 Spectral functions at Integer Filling
II.5.1 Half-Filling ($N_{e}=2$)

### v2

@@ II.2.2 Out-of-Plane Dimerization as Origin of Flat Band Gapping
Having established the in-plane breathing mode distortion and its accompanying trimerization as the mechanism producing the isolated flat bands, we are now ready to study their evolution throughout the Nb3X8 family. In Fig. 4, we show the bulk DFT band structures in the LT phases. As expected, we find two relatively flat bands around the Fermi level in all cases. In the case of Nb3F8, both bands cross the Fermi level, while for the others, we consistently find one fully occupied and one completely unoccupied band. For in-plane momenta ($k_{z}=0$), the two bands are clearly split apart, while for the out-of-plane momentum direction $\Gamma-A$, we observe a stronger dispersion, and the gap nearly closes. As will become clear from the subsequent Wannierization discussed below, this $k_{z}$ behaviour is a result of the interlayer dimerization of trimers within the strongly hybridized bilayers in the LT structure. Thus, in $k_{z}$, the low-energetic electronic structure is reminiscent of the behaviour of the Su-Schrieffer-Heeger (SSH) model [47], which consequently yields a gap in this direction, due to the mismatch of the alternating out-of-plane hopping matrix elements in the LT phase. From Cl to Br and I, we already see the first clear trend in the material class: the in-plane gap between these two bands steadily increases. In the HT phase, this out-of-plane dimerization is strongly suppressed yielding negligible gaps, rendering the low-energy band structures of all HT structures qualitatively similar to the LT Nb3F8* one.
II.2.3 Molecular Orbital Basis
@@
Figure 5: Wannier Molecular Orbitals of Nb3Br8. $\psi_{1}$ and $\psi_{2}$, along with their bonding (++) and anti-bonding (-) combinations, $\psi^{(\pm)}=(\psi_{1}\pm\psi_{2})/\sqrt{2}$.
@@
To prepare for the derivation of material-specific many-body Hamiltonians using ab initio downfolding and to further analyze the low-energetic band structure, we construct maximally localized Wannier functions for the two low-energy bands. In all four cases, we find that the bands can be perfectly Wannier-interpolated within a basis spanned by two molecular orbitals $\psi_{1}$ and $\psi_{2}$, which are centered on trimers of adjacent layers and which are directly above each other, see Fig. 5. As shown in Refs. [24, 21], these molecular orbitals arise from a direct linear combination of three Nb $d$ orbitals from the local trimers. The original atomistic multi-orbital basis has, however, very little influence due to the large energy separation between the trimer orbital and all other states, which are therefore either fully occupied or empty. In this molecular orbital basis, we find similar contributions of both orbitals to both low-energy bands. However, upon diagonalization of the single-particle Wannier Hamiltonian, we find that the lower (occupied) band is well described by a bonding orbital $\psi^{(+)}$ formed by the direct sum of the trimer-centered molecular orbitals $\psi^{(+)}=(\psi_{1}+\psi_{2})/\sqrt{2}$, while the upper (unoccupied) band is well described by the corresponding anti-bonding orbital $\psi^{(-)}=(\psi_{1}-\psi_{2})/\sqrt{2}$, as illustrated in Fig. 5. For the subsequent analysis of correlation effects this molecular orbital basis is most efficient, as it compresses six atomistic basis orbitals to just two molecular ones. Furthermore, it will be helpful to switch back and forth between the molecular orbital basis ($\psi_{i}$) and the bonding/anti-bonding basis ($\psi^{(\pm)}$), as the self-energy in the LT phase becomes diagonal in the bonding/anti-bonding basis.
@@
Upon projecting the Kohn-Sham states to the $\psi^{(\pm)}$ basis, we see that this interpretation holds especially for in-plane momenta. For out-of-plane $k_{z}$ momenta, however, we observe that $\psi^{(\pm)}$ states mix, which is a result of the small, but finite wavefunction overlap between the dimerized layers. This admixture thus verifies the SSH-like behaviour in the $k_{z}$ direction and is consistent with Ref. [48] within which Nb3Br8 and Nb3I8 are classified as obstructed atomic insulators  [^1] See https://www.topologicalquantumchemistry.fr/#/detail/25767 and https://www.topologicalquantumchemistry.fr/#/detail/421609.. The latter are characterized by wave-function centers that are not located at a atomic position, which holds in our case for the bonding combination of the trimer orbitals, cf. $\psi^{(+)}$ in Fig. 5. Although the Cl, Br, and I compounds all share the classification as obstructed atomic insulators, the details and specifically the size of the DFT gap change from material to material, as reflected in the Wannier hopping parameters in Tab. 1 and as discussed in more detail in the following section.
@@
II.3 Downfolded Many-Body Hamiltonians from Ab Initio Constrained Random Phase Approximation Calculations
Table 1: Structural and Ab Initio Downfolded Many-Body Hamiltonian Parameters for LT Nb3X8. In-plane ($a$) and out-of-plane ($c$) lattice parameters, molecular orbital spread ($\Omega$), nearest neighbor in-plane ($t_{\parallel}$) and out-of-plane ($t_{\perp}$) hopping parameters and corresponding cRPA screened Coulomb ($U_{\parallel}$ and $U_{\perp}$) interactions, on-site bare ($V_{0}$) and screened ($U_{0}$) Coulomb interactions including their ratio $\varepsilon_{\text{eff}}=V_{0}/U_{0}$.
X $a$ $c$ $\Omega$ $t_{\parallel}$ $t_{\perp}^{\text{s}}$ $t_{\perp}^{\text{w}}$ $V_{0}$ $U_{0}$ $\varepsilon_{\text{eff}}$ $U_{\parallel}$ $U_{\perp}^{\text{s}}$ $U_{\perp}^{\text{w}}$
F Bulk 5.74 31.16 4.12 34.3 -4.9 -6.5 7486.4 2590.5 2.89 804.5 714.6 572.5
BL 33.8 -5.0 7496.2 3988.8 1.88 1977.2 1987.7
ML 33.6 7497.2 4021.8 1.86 2016.7
Cl Bulk 6.75 36.75 7.08 20.1 -136.0 -16.1 6126.4 1451.4 4.22 475.6 400.1 313.5
BL 18.5 -136.2 6150.0 2697.6 2.28 1598.5 1570.9
ML 18.6 6180.3 2771.3 2.23 1670.1
Br Bulk 7.08 40.84 8.54 5.9 -169.4 -20.4 5787.4 1186.6 4.88 381.5 342.0 262.4
BL 3.4 -169.2 5819.4 2396.0 2.43 1483.8 1482.3
ML 4.5 5887.7 2475.3 2.38 1554.3
I Bulk 7.70 43.24 13.15 -8.3 -218.2 -24.6 5052.7 787.0 6.42 258.3 258.5 183.8
BL -13.2 -218.4 5101.8 1928.7 2.65 1323.3 1349.0
ML -11.5 5321.1 2052.4 2.59 1397.3
@@ II.3 Downfolded Many-Body Hamiltonians from Ab Initio Constrained Random Phase Approximation Calculations
for all compounds. Here $c_{i}^{\dagger}$ and $c_{j}$ are electron creation and annihilation operators in molecular orbitals at sites $i$ and $j$, respectively, $t_{ij}$ are single-particle (hopping) and $U_{ijkl}$ partially screened two-particle (Coulomb) matrix elements. The latter are evaluated within the constrained random phase approximation (cRPA) [50], which accounts for all RPA-level screening processes except those taking place within the molecular orbitals $\psi_{i}$. In this way, all intra- and interlayer screening properties are included. As a first-principles method, the cRPA therefore captures how screening differs across the class of materials, both due to changes in the chemical composition as well as due to differences in the screening environment between monolayers, bilayers, and bulk structures. This way, our material-specific localized molecular orbitals render adequately the different local chemical environments in each compound and in each stacking, while the screening captures global crystal effects including polarizations from higher and deeper lying electronic states.
The cRPA scheme gives access to the full tensor $U_{ijkl}$. In Tab. 1, we show a subset of these matrix elements focusing on the local and nearest-neighbor density-density Coulomb ($U_{ijji}$) as well as on intra- and inter-layer hopping matrix elements. For a complete overview and details about the calculations, see Section V.1.
We start by discussing the molecular orbital spreads ($\Omega$) of the bulk structures, which we find to increase from F to I. This trend follows the lattice constants, which increase in the in- and out-of-plane direction from F to I. This shows how the molecular orbitals and thus the local orbital basis reflects the chemical changes in the structure. The tightest-bound lattice (Nb3F8*) hosts highly localized molecular orbitals, while the weakest-bound lattice (Nb3I8) has rather delocalized molecular orbitals.
@@
These trends are further reflected in the hopping matrix elements. Overall, we find that the in-plane nearest-neighbor hopping ($t_{\parallel}$) shows a decreasing trend from F to I. Its magnitude is relatively small in all compounds, which reflects the flat in-plane (constant $k_{z}$) dispersion. The strong out-of-plane nearest-neighbor hopping ($t^{(s)}_{\perp}$), in turn, grows significantly in magnitude from F to I, which is in line with a decrease in the orbital spread. The “weak link” in the out-of-plane directions ($t^{(w)}_{\perp}$) is in all compounds about an order of magnitude smaller than $t^{(s)}_{\perp}$ and accordingly increases in magnitude slightly from F to I. This reflects the increasing bilayer hybridization between the inter-layer dimerized trimers, leading to the growth of the bonding/anti-bonding single particle splitting, as observed around $\Gamma$ in Fig. 4. In Nb3F8*, the in- and out-of-plane hoppings are not very different, such that there is no opening of a full dimerization gap throughout the Brillouin zone and only a small splitting at $\Gamma$ is observed. To analyze the gap at the high symmetry point $A$, we need to take into account that there are three adjacent layer neighbors in the “weak-link” directions, such that the SSH gap of a nearest-neighbor model would be given by $\tau_{\perp}=2\left|t^{(s)}_{\perp}-\sqrt{3}t^{(w)}_{\perp}\right|$. This increases from about 216 to 351  meV going from Cl to I, indicating an increase of the dimerization strength (in Nb3F8*, it is only 44\,meV). The DFT gaps at $A$ (where the SSH gap opens) are, however of similar size in Nb3Cl8 and Nb3Br8. We understand this as a result of long-range hoppings suppressing the dimerization, particularly hopping to next nearest and next-next-nearest neighbouring trimers in the adjacent van der Waals layer. These hoppings become slightly more pronounced from Cl to I.
The most important interaction parameter is the partially screened local density-density interaction $U_{0}$ (Hubbard’s $U$), which depends on the size of the orbital and on the screening. We can disentangle these effects by also looking at the bare local density-density Coulomb interaction $V_{0}$, which depends only on the size of the orbital. It is largest in Nb3F8* and smallest in Nb3I8, inversely following the molecular orbital spread. The impact of screening can then be quantified by the ratio $\varepsilon_{\text{eff}}=V_{0}/U_{0}$, which is seen to increase drastically from F to I. This screening parameter is approximately inversely proportional to the “background” electronic gap (i.e., the gap between the filled and occupied bands above and below the flat bands). Thus, the screened local density-density Coulomb interactions are not just suppressed by the enhanced molecular orbital spread, but also by stronger screening going from F to I.
@@
Next to this, our calculations show that the partially screened Coulomb interactions are dominated by density-density interactions. This is a result of having just a single molecular orbital being centered in each of the adjacent, strongly hybridized monolayers. Non-density-density interactions, such as Hund’s exchange interactions, are thus non-local, which significantly suppresses their strengths (to a few meV). Due to reduced screening in the layered vdW structures, the density-density Coulomb interactions are also non-local and show a significant long-ranged tail. The latter is visible in Tab. 1 by comparing the local interactions $U_{0}$ to the nearest-neighbor ones in both in- and out-of-plane direction. In contrast to the out-of-plane hopping matrix elements, we do not see a significant bilayer hybridization effect in the non-local density-density Coulomb interactions. Comparing $U^{(s)}_{1,\perp}$ and $U^{(w)}_{1,\perp}$, we find only small differences, which can be explained by the slightly different distances between the molecular orbital centers in these two directions.
Comparing the bulk parameter with the mono- and bilayer ones, we do not see sizable differences in the hopping parameter and only mild enhancements of the bare interactions, but strong enhancements of the screened Coulomb matrix elements. The latter is mostly driven by the reduced screening in the mono- and bilayer limit as compared to the bulk structures, as clearly visible from the reduced $\varepsilon_{\text{eff}}$ in these structures. Thus, while the bulk and bilayer structures behave rather similar in terms of their non-interacting electronic structure, they could be vastly different as a result of different screening environments.
@@
Taken together, our ab initio downfolding calculations show that the local Coulomb interactions $U_{0}$ are very large in comparison to in-plane hopping terms with $U_{0}/t_{\parallel}$ ratios always larger than $80$. In comparison to the strong out-of-plane hopping matrix elements, we find $U_{0}/t_{\perp}$ ratios that decrease from about $12$ in Nb3Cl8 to around $4$ in Nb3I8. These ratios point towards possible strong Coulomb-interaction induced correlation effects, which might change from system to system. To investigate this and quantify the correlation strengths, we proceed with solving the generalized Hubbard models using cluster dynamical mean-field theory.
@@
II.4 Generalized Phase Diagram from Cluster Dynamical Mean Field Theory at Half-Filling
II.4.1 Self-Energies
@@
Figure 6: Nb3X8 Matsubara Self-Energies. (a) Real and (b) imaginary parts of the self energies in Matsubara frequencies in the bonding/anti-bonding basis. Markers represent Hubbard-I results, lines represent cluster DMFT results.
@@
Using our cRPA-based downfolded molecular-orbital models, we calculate the self-energies at half-filling within dynamical mean-field theory (DMFT) and its Hubbard-I approximation [51, 52]. Specifically, we define the impurity problem as the two inter-layer dimerized molecular (trimer) orbitals within each bilayer and include the local $U_{0}$, the strong inter-layer hybridization $t^{(s)}_{\perp}$ as well as all inter-layer Coulomb interaction elements $U_{ijkl}$ (among which $U^{(s)}_{\perp}$ is by far the largest, see Tab. 1). Accordingly, we deal with a cluster DMFT problem [53, 54, 19]. Fig. 6 shows the resulting cluster DMFT Matsubara frequency self-energies in the diagonal $\psi^{(\pm)}$ basis for all compounds calculated using a CT-HYB quantum Monte Carlo solver at $\beta=100\,$eV-1 (see V.1 for all details). In this basis at half-filling, the imaginary parts of the self-energy are the same for both channels ($\mathrm{Im}\Sigma_{(+)}=\mathrm{Im}\Sigma_{(-)}$), while the real parts differ only in their signs ($\mathrm{Re}\Sigma_{(+)}=-\mathrm{Re}\Sigma_{(-)}$) in the case of Cl, Br, and I. For Nb3F8* the bonding and anti-bonding self-energy lie on top of each other in Fig. 6.
@@
The real part, as depicted in Fig. 6 (a), shows that correlation effects enhance the bonding/anti-bonding splitting, similar to the Stoner enhancement of the exchange splitting in magnetic systems. In I, the real part is almost frequency-independent, while it becomes strongly frequency-dependent moving towards Cl. This is a sign of the increasing importance of dynamic correlation effects, which necessitate (cluster) DMFT calculations. In the imaginary part depicted in Fig 6 (b), we again find a clear trend: the overall amplitudes of the self-energy gradually increase from I to Br, Cl, and F, which is another sign of the increasing degree of correlation. Specifically, we see that the imaginary part of the self-energy of Nb3I8 is barely changing with frequency, while it clearly shows retardation effects in Nb3Cl8 and Nb3Br8. Finally, Nb3F8* shows a divergent imaginary self-energy for small Matsubara frequencies.
@@
II.4.2 Phase Classification Based on Local Molecular Dimer Properties
Figure 7: Local Dimer Analysis. (a) shows the charge neutral excitation spectrum. (b) and (c) show the temperature dependence of expectation values of $\braket{n_{0,\uparrow}n_{0,\downarrow}}$ double occupation and inter-layer spin alignment $\braket{n_{0,\uparrow}n_{1,\uparrow}}$. These can be interpreted as measures for the thermal occupation of the first charge and spin excitation respectively. This can be seen by clear transitions in behavior at the characteristic temperatures $T_{c}=E_{exc}/k_{b}$, plotted as dotted/dashed lines. The black line is the temperature at which we perform our DMFT study $\beta=100eV^{-1}$.
To further substantiate this classification of LT bulk Nb3X8 into Mott, strongly and weakly correlated insulators, we proceed with analyzing the “local impurity”, i.e. the dimer formed by the hybridized molecular orbitals from adjacent monolayers, in more detail. This is similar to the analysis by Zhang et al. [22]. However, we use here our material-specific model parameters and, importantly, the full Coulomb tensor including non-local interactions between the molecular orbitals sites (which is important for charge-neutral excitations [55]) to not only calculate $T=0$ properties, but also finite temperature expectation values.